# Explaining the Behavior of Black-Box Prediction Algorithms with Causal Learning

### v5

@@ Appendix B Appendix: Simulation Details And Implementation
We measure model performance on a validation dataset of $500$ images, and our final network achieves the following accuracy on the validation dataset. A $100\%$ accuracy in predicting circles, $99.60\%$ for predicting horizontal bars, $97.80\%$ accuracy for predicting triangles and $100\%$ accuracy in predicting vertical bars. Our logistic regression model is using the sklearn library, with no penalty. The learned intercept is $-1.82530524$, and the learned coefficients are $0.09042825$, $1.56043942$, $0.07410451$, and $1.27937188$ for circle, horizontal bar, triangle and vertical bar respectively. This full black-box model achieves $77.89\%$ accuracy on the dataset generated according to the DGP outlined in the simulation section.
For the FCI algorithm, we use the implementation in the TETRAD package, with the Chi-square independence test at a significance level of $0.005$. Additionally, we impose the background knowledge that $\widehat{Y}$ must be a descendant of the interpretable features.
Appendix C Appendix: Preprocessing for Bird Dataset
@@
C.1 Bird Classification Grouping

### v6

@@ Appendix B Appendix: Simulation Details And Implementation
We measure model performance on a validation dataset of $500$ images, and our final network achieves the following accuracy on the validation dataset. A $100\%$ accuracy in predicting circles, $99.60\%$ for predicting horizontal bars, $97.80\%$ accuracy for predicting triangles and $100\%$ accuracy in predicting vertical bars. Our logistic regression model is using the sklearn library, with no penalty. The learned intercept is 1.82530524-1.82530524, and the learned coefficients are $0.09042825$, $1.56043942$, $0.07410451$, and $1.27937188$ for circle, horizontal bar, triangle and vertical bar respectively. This full black-box model achieves $77.89\%$ accuracy on the dataset generated according to the DGP outlined in the simulation section.
For the FCI algorithm, we use the implementation in the TETRAD package, with the Chi-square independence test. Additionally, we impose the background knowledge that $\widehat{Y}$ must be a descendant of the interpretable features.
Appendix C Appendix: sensitivity analysis with hybrid causal discovery algorithm
As an alternative to FCI, we also apply a hybrid procedure called GRaSP-FCI to the data on chest X-ray and bird images. There is no fully greedy score-based procedure for learning PAGs analogous to the GES (Greedy Equivalence Search) algorithm for learning CPDAGs (Chickering, 2002). However, a two-step hybrid procedure has been proposed: first learn a CPDAG by using a greedy score based procedure, and then ignore edge directions and post-process the resulting undirected graph with some additional independence tests and FCI-orientation rules to produce a PAG (Ogarrio et al., 2016). Details underlying this hybrid approach and a proof of its correctness are described in Ogarrio et al. (2016). Though the authors use GES as the greedy score-based algorithm in that paper, more recently it has become possible to substitute the GRaSP (Greedy Relaxations of Sparsest Permutation) algorithm for learning a CPDAG (Lam et al., 2022) in the first stage of this two-step procedure; the result is called GRaSP-FCI. This algorithm is implemented in the TETRAD package and thus straightforward to run on the same data with the same pipeline (including the background knowledge described above). We use the BDeu score and Chi-square test of independence, with all other settings the same as for FCI. Results are summarized in Figure 15. In the X-ray image analysis, the ordering of features by edge stability is the same for GRaSP-FCI as FCI. The ordering for the bird image analysis has substantive differences: the edge stabilities are overall lower, and the top variables in the ordering are Crown Color and Size.
Figure 15: Results for two experiments using GRaSP-FCI. Potential causal determinants of (a) bird classifier output from bird images and (b) pneumonia classifier output from X-ray images.
Appendix D Appendix: Preprocessing for Bird Dataset
@@
D.1 Bird Classification Grouping